# TIF-UNIMI-2024-12 Asymptotic Freedom in Parton Language: the Birth

### v1

@@ 1 Parisi and QCD
These contributions can be broadly subdivided into three partly overlapping epochs. A first epoch, from 1970 until about 1973, during which the main issue was understanding the meaning of the parton model, explaining its phenomenological success, and understanding its possible theoretical underpinnings and its implications. In other words, understanding Bjorken scaling: the surprising discovery that structure functions, the form factors that describe the structure of the proton as probed by the deep-inelastic scattering of electrons (DIS) are approximately scale dependent. A second epoch, ranging from 1972 to about 1976, in which the nature of QCD as a quantum field theory starts emerging, and the parton model is understood as a limit of perturbative QCD — culminating with the seminal paper [7] that gives its name to this contribution. In other words, understanding scaling violations: the weak logarithmic scale dependence of hadronic structure. And a last epoch, from 1976 to the end of the decade, in which the foundations of perturbative QCD are laid, by introducing some of its basic ideas and developing its fundamental tools.
@@ 2 Scaling and the parton model
i.e. the ratio of hadron to muon-antimuon production cross-sections. The well-known leading-order result , where are the electric charges of partons, i.e. the constituent of the proton as probed in the perturbative domain, is given as Eq. (1) of this paper.
@@
for the proton-neutron difference of the structure functions that parametrize the deep-inelastic cross-section. Amusingly, in more recent textbooks (see e.g. Ref. [13]) this result is derived using naive parton model argument, and referred to as the Gottfried [14] sum rule — though it is nowhere to be found in Gottfried’s paper [14]. The discovery of its experimental violation in the early nineties arguably opened up the modern era of understanding of the parton substructure of the nucleon [15].
@@
Before turning to the line of thought that led to this celebrated result, it is worth spending a few words to discuss some studies that were part of investigations of QED, rather than QCD, but are precursors of techniques that are part of the standard QCD toolbox, and indeed some of which played an important role in subsequent QCD studies of Giorgio’s. The first [17] is a discussion of the process in QED in the hard Bremsstrahlung limit, in which the photon energy is large. This paper is based on the use of the Weizsäcker-Williams, or quasi-real photon approximation, in which the photon emission rate is computed in terms of a universal probability for the electron to split into a photon. This approximation will later play an important role in the derivation of evolution equations for parton distributions, namely, the Altarelli-Parisi equations, that I will discuss in Sect. 3 below. In this paper, the result found using this approximation is compared to the exact result. The approximation holds in the limit of collinear photon emission, but it is shown to provide an acceptable approximation for surprisingly large angles — a fact that, with hindsight, explains the unexpected phenomenological success of the Altarelli-Parisi equation, and consequently of Bjorken scaling, even at the relative low scales that were experimentally probed at that time, a phenomenon that was later called precocious scaling.
The second is the companion paper [18], in which the exact computation for this QED process is presented. In this work it is observed that, when computing Feynman diagrams with several particles in the final state, the use of helicity amplitudes instead of traces is highly advantageous. The computation is then performed using helicity amplitude methods, now standard, but only popularized by others many years later [19]. The computational technique itself is in fact the main point of the paper, whose stated goal in its abstract is to “propose a general method to evaluate Feynman graphs” which “avoids the calculation of very long traces”.
@@ 3 Scaling violations
It is then observed that if one tries to apply the same reasoning to QCD, namely to the structure of the proton, rather than the electron, there is a fundamental difference, namely, that in QED the vacuum behaves as a dielectric, and consequently the electric charge is screened at large distances, and increases at shorter distances (an argument that is now textbook [9]). In QCD the opposite happens: strongly interacting matter is an “enantion” (loosely translatable as a “contrarian”), it anti-screens, and consequently the charge decreases at shorter distances — asymptotic freedom. Combining these two physical pictures, scale dependence from radiation, and asymptotic freedom from anti-screening, a simple physical picture of the scaling laws of proton structure in QCD emerges. Hence, the key insight in these lectures is to interpret the scaling laws obeyed by deep-inelastic structure function as a consequence of the scaling law of the number densities of partons — the proton’s constituents. This then provides a clear physical picture of scale dependence from evolution equations satisfied by structure functions as a showering process, as realized in modern Monte Carlo parton shower codes [28]. The solution to these evolution equations corresponds to summing a class of Feynman diagrams that had been considered previously by Gribov and Lipatov, who had studied deep-inelastic scattering in a theory with massive vector mesons a few years earlier [29]. However, the derivation presented here is completely independent, and in fact the physical picture of evolution equations is missing in the Gribov-Lipatov approach.
All the pieces of the puzzle come together in the subsequent celebrated, seminal Altarelli-Parisi paper [7] (of which the Moriond lectures [27] are presented as a “preliminary, less complete version”). This is of course one of the most widely cited papers in high-energy physics. In this paper the focus shifts from the deep-inelastic structure function, to the structure of the proton itself. Indeed, the first insight is that the scale dependence of structure function is a reflection of an underlying scale dependence of the structure of the proton, as reflected by parton distribution functions, that provide number densities for the constituents of the proton. The second insight is that the previous mathematical observation [20, 23], that scaling laws can be obtained by solving an integro-differential equations and taking the Mellin transform of the result, can be combined with the physical idea that the proton constituents can be viewed as composite systems, an idea previously suggested by Kogut and Susskind [30].
@@
The derivation of these equations presented in Ref. [7] is arguably rather more transparent than the majority of current available textbook treatments. However, their significance goes well beyond the technical aspects of their derivation. In fact, the fundamental meaning of this work is sometimes not fully appreciated. Indeed, this work is sometimes described [9] as an independent re-derivation of the previous argument by Gribov and Lipatov [29], that shows how to derive evolution equations for deep-inelastic structure functions themselves, rather than their moments. Whereas of course the Altarelli-Parisi paper does contain, as it is stated in its introduction “an alternative derivation of all results of current interest for the behavior of deep inelastic structure functions” (alternative to the derivation based on the scale dependence of matrix elements of leading-twist operators, that is), this is not the main point. Rather, the point, as stated in the title of the paper, is to express quantum-field theoretical results of QCD in parton language. In fact, this work brings to completion the research program and line of thought that started in the early papers that asked the questions: what is the nature of partons? and why does the parton model work?
Indeed, even in modern textbook treatments [10] a parton model language and parton-based approach to QCD factorization is often taken for granted: it is assumed that hard processes can be described by starting with the parton model, and computing perturbative corrections to it. Yet, the theoretical justification for this is unclear — it is unclear why the simple assumptions of the parton model are a reasonable starting point111See e.g. Sect. 9.11 of Ref. [33] for a modern critique of a naive parton-based approach.. What the Altarelli-Parisi paper shows, is that, for deep-inelastic structure functions, a parton-based computation is mathematically equivalent to the result found starting from the Wilson expansion, and then solving renormalization-group equations for Wilson coefficients. Indeed, a terse comment in the conclusion explains “the reasons for the success of this simple method” — the renormalization group equations. So what the Altarelli-Parisi paper amounts to, is a rigorous proof of QCD factorization, albeit in the limited case of deep-inelastic scattering. The paper is at the basis of the modern description of the structure of the proton in terms of parton distributions (see Ref. [34] and ref. therein) but its true significance is in establishing perturbative factorization.
@@ 4 The birth of QCD phenomenology
With QCD factorization on a solid field-theoretic basis, at least for deep-inelastic scattering, the leap to full-fledged QCD phenomenology was the logical development. This was laid out in a series of papers, submitted within the short span of two years, which set the basis of QCD calculations for hadronic processes [35, 36], the definition of jet substructure observables [37, 2], the determination of parton distributions [38], the scale dependence of fragmentation functions [39], transverse momentum resummation [40], and threshold resummation [41]. Many of these contributions anticipate ideas of methods that have marked the development of QCD phenomenology, sometimes several decades later.
Some of these papers had a large impact, and shaped subsequent developments, though some were only recognized at a later stage, and some not at all. In particular, Ref. [35] presents a first computation of the dimuon transverse momentum distribution in Drell-Yan production. The celebrated Collins-Soper-Sterman (CSS) work [42], that later set the modern standard for the perturbative calculation for this process, refers to Ref. [35] as the foundation of the whole field: “with the advent of QCD factorization one was able to write the DY cross-section at measured ”. Interestingly, the subsequent paper Ref. [36] considers the effect of taking into account the transverse momentum of partons, a topic that has only become fashionable in this century, with studies of transverse-momentum dependent parton distributions (TMDs).
@@ 5 The work and its context
I am grateful to the organizers of the seminar series “The interdisciplinary contribution of Giorgio Parisi to theoretical physics”, that took place at Sapienza University of Rome in 2022-2023, for inviting me to give a talk in the series, and for subsequently taking the initiative of collecting in a volume the write-ups of all talks. Unlike most of the other contributors, I did not have the privilege to collaborate with Giorgio directly. Somehow, despite my longtime association with my late teacher, mentor and friend Guido Altarelli, I never managed to discuss his joint work with Parisi with him. Moreover, Parisi’s QCD work was done at a time when I was a school kid, so I could not witness its development first-hand. I consequently had to approach it with the attitude of a historian (which I am not), and to some extent of a teacher (which I certainly am). I thus have the obligation of apologizing in advance for any misunderstanding of which I might well have been responsible. Among the organizers, I am especially grateful to Marco Bonvini and Maria Chiara Angelini for comments and criticism on a preliminary version of this manuscript. I would finally like to thank Giorgio for graciously sitting through my talk and sharing with me some comments, thoughts and anecdotes.

### v2

@@ 1 Parisi and QCD
These contributions can be broadly subdivided into three partly overlapping epochs. A first epoch, from 1970 until about 1973, during which the main issue was understanding the meaning of the parton model, explaining its phenomenological success, and understanding its possible theoretical underpinnings and its implications. In other words, understanding Bjorken scaling: the surprising discovery that structure functions, the form factors that describe the structure of the proton as probed by the deep-inelastic scattering of electrons (DIS) are approximately scale independent. A second epoch, ranging from 1972 to about 1976, in which the nature of QCD as a quantum field theory starts emerging, and the parton model is understood as a limit of perturbative QCD — culminating with the seminal paper [7] that gives its name to the present paper. In other words, understanding scaling violations: the weak logarithmic scale dependence of hadronic structure. And a last epoch, from 1976 to the end of the decade, in which the foundations of perturbative QCD are laid, by introducing some of its basic ideas and developing its fundamental tools.
@@ 2 Scaling and the parton model
i.e. the ratio of hadron to muon-antimuon production cross-sections. Equation (1) of this paper states the now well-known leading-order result: , where are the electric charges of partons, i.e. the constituents of the proton as probed in the perturbative domain, and the sum runs over all constituents, and thus can be used to establish both the nature and the total number of constituents — and thus eventually provided evidence for the existence of three quark colors.
@@
for the proton-neutron difference of the structure functions that parametrize the deep-inelastic cross-section. Amusingly, in more recent textbooks (see e.g. Ref. [13]) this result is derived using naive parton model arguments, and referred to as the Gottfried [14] sum rule — though it is nowhere to be found in Gottfried’s paper [14]. The discovery of its experimental violation in the early nineties arguably opened up the modern era of understanding of the parton substructure of the nucleon [15].
@@
Before turning to the line of thought that led to this celebrated result, it is worth spending a few words to discuss some studies that were part of investigations of QED, rather than QCD, but are precursors of techniques that are now part of the standard QCD toolbox, and indeed some of which played an important role in subsequent QCD studies of Giorgio’s. The first [17] is a discussion of the process in QED in the hard Bremsstrahlung limit, in which the photon energy is large. This paper is based on the use of the Weizsäcker-Williams, or quasi-real photon approximation, in which the photon emission rate is computed in terms of a universal probability for the electron to split into a photon. This approximation will later play an important role in the derivation of evolution equations for parton distributions, namely, the Altarelli-Parisi equations, that I will discuss in Sect. 3 below. In this paper, the result found using this approximation is compared to the exact result. The approximation holds in the limit of collinear photon emission, but it is shown to provide an acceptable approximation for surprisingly large angles. With hindsight, this explains the unexpected phenomenological success of the Altarelli-Parisi equation, and consequently of Bjorken scaling, even at the relative low scales that were experimentally probed at that time, a phenomenon that was later called precocious scaling.
The second is the companion paper [18], in which the exact computation for this QED process is presented. In this work it is observed that, when computing Feynman diagrams with several particles in the final state, it is highly advantageous to use, instead of Dirac traces, the helicity amplitude method, previously introduced in Ref. [19]. The computation is then performed using this method, now standard, but only popularized by others many years later [20]. The computational technique itself is in fact the main point of the paper, whose stated goal in its abstract is to “propose a general method to evaluate Feynman graphs” which “avoids the calculation of very long traces”.
@@ 3 Scaling violations
It is then observed that if one tries to apply the same reasoning to QCD, namely to the structure of the proton, rather than the electron, there is a fundamental difference, namely, that in QED the vacuum behaves as a dielectric, and consequently the electric charge is screened at large distances, and increases at shorter distances (an argument that is now textbook [9]). In QCD the opposite happens: strongly interacting matter is, in the words of Ref. [28], an “enantion”, loosely translatable as a “contrarian”: it anti-screens, and consequently the charge decreases at shorter distances — asymptotic freedom. Combining these two physical pictures, scale dependence from radiation, and asymptotic freedom from anti-screening, a simple physical picture of the scaling laws of proton structure in QCD emerges. Hence, the key insight in these lectures is to interpret the scaling laws obeyed by deep-inelastic structure function as a consequence of the scaling law of the number densities of partons — the proton’s constituents. This then provides a clear physical picture of scale dependence from evolution equations satisfied by structure functions as a showering process, as realized in modern Monte Carlo parton shower codes [29]. The solution to these evolution equations corresponds to summing a class of Feynman diagrams that had been considered previously by Gribov and Lipatov, who had studied deep-inelastic scattering in a theory with massive vector mesons a few years earlier [30]. However, the derivation presented here is completely independent, and in fact the physical picture of evolution equations is missing in the Gribov-Lipatov approach.
All the pieces of the puzzle come together in the subsequent celebrated, seminal Altarelli-Parisi paper [7] (of which the Moriond lectures [28] are presented as a “preliminary, less complete version”). This is of course one of the most widely cited papers in high-energy physics. In this paper the focus shifts from the deep-inelastic structure function, to the structure of the proton itself. Indeed, the first insight is that the scale dependence of structure function is a consequence of an underlying scale dependence of the structure of the proton, as reflected by parton distribution functions, that provide number densities for the constituents of the proton. The second insight is that the previous mathematical observation [21, 24], that scaling laws can be obtained by solving an integro-differential equations and taking the Mellin transform of the result, can be combined with the physical idea that the proton constituents can be viewed as composite systems, an idea previously suggested by Kogut and Susskind [31].
@@
The derivation of these equations presented in Ref. [7] is arguably rather more transparent than the majority of current available textbook treatments. However, their significance goes well beyond the technical aspects of their derivation. In fact, the fundamental meaning of this paper is sometimes not fully appreciated. Indeed, this work is sometimes described [9] as an independent re-derivation of the previous argument by Gribov and Lipatov [30], that shows how to obtain evolution equations for deep-inelastic structure functions themselves, rather than their moments. Whereas of course the Altarelli-Parisi paper does contain, as it is stated in its introduction “an alternative derivation of all results of current interest for the behavior of deep inelastic structure functions” (alternative to the derivation based on the scale dependence of matrix elements of leading-twist operators, that is), this is not the main point. Rather, the point, as stated in the title of the paper, is to express quantum-field theoretical results of QCD in parton language. In fact, this work brings to completion the research program and line of thought that started in the early papers that asked the questions: what is the nature of partons? and why does the parton model work?
Indeed, even in modern textbook treatments [10] a parton model language and parton-based approach to QCD factorization is often taken for granted: it is assumed that hard processes can be described by starting with the parton model, and computing perturbative corrections to it. Yet, the theoretical justification for this is unclear — it is unclear why the simple assumptions of the parton model are a reasonable starting point111See e.g. Sect. 9.11 of Ref. [34] for a modern critique of a naive parton-based approach.. What the Altarelli-Parisi paper shows, is that, for deep-inelastic structure functions, a parton-based computation is mathematically equivalent to the result found starting from the Wilson expansion, and then solving renormalization-group equations for Wilson coefficients. Indeed, a terse comment in the conclusion explains “the reasons for the success of this simple method” — the renormalization group equations. So what the Altarelli-Parisi paper amounts to, is a rigorous proof of QCD factorization, albeit in the limited case of deep-inelastic scattering. The paper is at the basis of the modern description of the structure of the proton in terms of parton distributions (see e.g. Ref. [35] and ref. therein) but its true significance is in establishing perturbative factorization.
@@ 4 The birth of QCD phenomenology
With QCD factorization on a solid field-theoretic basis, at least for deep-inelastic scattering, the leap to full-fledged QCD phenomenology was the logical development. This was laid out in a series of papers, submitted within the short span of two years, which set the basis of QCD calculations for hadronic processes [36, 37], the definition of jet substructure observables [38, 2], the determination of parton distributions [39], the scale dependence of fragmentation functions [40], transverse momentum resummation [41], and threshold resummation [42]. Many of these contributions anticipate ideas and methods that have marked the development of QCD phenomenology, sometimes several decades later.
Some of these papers had a large impact, and shaped subsequent developments, though some were only recognized at a later stage, and some not at all. In particular, Ref. [36] presents a first computation of the dimuon transverse momentum distribution in Drell-Yan production. The celebrated Collins-Soper-Sterman (CSS) work [43], that later set the modern standard for the perturbative calculation for this process, refers to Ref. [36] as the foundation of the whole field: “with the advent of QCD factorization one was able to write the DY cross-section at measured ”. Interestingly, the subsequent paper Ref. [37] considers the effect of taking into account the transverse momentum of partons, a topic that has only become fashionable in this century, with studies of transverse-momentum dependent parton distributions (TMDs [44]).
@@ 5 The work and its context
I am grateful to the organizers of the seminar series “The interdisciplinary contribution of Giorgio Parisi to theoretical physics”, that took place at Sapienza University of Rome in 2022-2023, for inviting me to give a talk in the series, and for subsequently taking the initiative of collecting in a volume the write-ups of all talks. Unlike most of the other contributors, I did not have the privilege of collaborating with Giorgio directly. Somehow, despite my longtime association with my late teacher, mentor and friend Guido Altarelli, I never managed to discuss his joint work with Parisi with him. Moreover, Parisi’s QCD work was done at a time when I was a school kid, so I could not witness its development first-hand. I consequently had to approach it with the attitude of a historian (which I am not), and to some extent of a teacher (which I certainly am). I thus have the obligation of apologizing in advance for any misunderstanding of which I might well have been responsible. Among the organizers, I am especially grateful to Marco Bonvini and Maria Chiara Angelini for comments and criticism on a preliminary version of this manuscript. I also thank Giovanni Stagnitto for spotting several typos. I would finally like to thank Giorgio for graciously sitting through my talk and sharing with me some comments, thoughts and anecdotes.